OpenStudy (anonymous):

find value of x and y (1+i) y ka squre + (6+i) = (2+i)x

sam (.sam.):

\[ (1+i) y ka^2 + (6+i) = (2+i)x \]?

OpenStudy (anonymous):

heh listen if u r not able to understnd can i attached the file and tell u whch qustion is neded to be done

sam (.sam.):

Yeah attach it

OpenStudy (anonymous):

r u able to see the qustionss

OpenStudy (anonymous):

i want answers of question no. 2 b and c part and q3 and q1. f part pllzzz help mee

OpenStudy (gorv):

((1+i)y)^2+(6+i)=(2+i)x (1+i)^2*y^2+6+i=2x+ix (1-1+2*i*1)*y^2+6+i=2x+ix i2y^2+6+i=2x+ix 6+i(2y^2+1)=2x+ix now compare real and imaginary part of both side 2x=6 and x=2y^2+1 x=3 3=2y^2+1 2=2y^2 y^2=1 y= +-1 ans (x,y)=(3,1) or (3,-1)

OpenStudy (anonymous):

which answer is this

OpenStudy (anonymous):

^ yeh symbol kiska hai

OpenStudy (gorv):

y^2 means y square

OpenStudy (gorv):

this is the answer of your question

OpenStudy (anonymous):

kk thanks can u solve more qustions that i have post

OpenStudy (gorv):

yeah tell me

OpenStudy (anonymous):

i have attached the qustions in this i need answer of q1.f) part and q2 c part and q3. answer plzzzzzz provide me

OpenStudy (anonymous):

anyoneeeeeeee thereeeeee